OpenStudy (bananas1234):

Find a number that is between 8/11 and 0.85 repeating 23/33 39/66 157/198 88/99

OpenStudy (bananas1234):

could you help @zepdrix ?

OpenStudy (bananas1234):

Could you help? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you able to convert 8/11 to decimal form?

OpenStudy (bananas1234):

0.72 repeating

OpenStudy (bananas1234):

so i have to find what between 0.72 and 0.85. right?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

convert each answer choice to a decimal number as well

OpenStudy (bananas1234):

88/99 = 0.88 repeating 39/66 = 0.5909 09 repeating 157/198 = 0.99 23/33 = 0.69 repeating

OpenStudy (bananas1234):

ok i did that

jimthompson5910 (jim_thompson5910):

which of those decimal values are between 0.727272727272... and 0.858585858585...

jimthompson5910 (jim_thompson5910):

157/198 = 0.99 is incorrect btw

OpenStudy (bananas1234):

0.69 repeating or 23/33

OpenStudy (bananas1234):

sorry i wasn't sure about that one

jimthompson5910 (jim_thompson5910):

157/198 = 0.7929292929292...

OpenStudy (bananas1234):

ok

jimthompson5910 (jim_thompson5910):

23/33 = 0.696969... 39/66 = 0.590909... 157/198 = 0.79292929... 88/99 = 0.88888... which of those decimal values are between 0.727272727272... and 0.858585858585...

OpenStudy (bananas1234):

The first one

OpenStudy (dan815):

u know how to find fractions that repeat in general?

OpenStudy (bananas1234):

no

OpenStudy (rational):

how can 0.69 be between 0.72 and 0.85 ?

OpenStudy (bananas1234):

oh never mind then.....it must be 0.7929292929.....

OpenStudy (rational):

You got it!

OpenStudy (bananas1234):

Awesome thanks everyone!!

OpenStudy (rational):

Dont you find it ridiculous to see a cute fraction like 23/33 turn into messy looking decimal form : 0.696969696969696969... ?

OpenStudy (dan815):

hmm wud we go back with geometric series

OpenStudy (bananas1234):

Yes!! i have a really hard time with fractions. They seem simple but they are kinda confusing to me anyway.

OpenStudy (dan815):